# Privacy-Preserving Neural Graph Databases

### 19 Jan 2024 v2

@@ Abstract.
In the era of big data and rapidly evolving information systems, efficient and accurate data retrieval has become increasingly crucial. Neural graph databases (NGDBs) have emerged as a powerful paradigm that combines the strengths of graph databases (graph DBs) and neural networks to enable efficient storage, retrieval, and analysis of graph-structured data. The usage of neural embedding storage and complex neural logical query answering provides NGDBs with generalization ability. When the graph is incomplete, by extracting latent patterns and representations, neural graph databases can fill gaps in the graph structure, revealing hidden relationships and enabling accurate query answering. Nevertheless, this capability comes with inherent trade-offs, as it introduces additional privacy risks to the database. Malicious attackers can infer more sensitive information in the database using well-designed combinatorial queries, such as by comparing the answer sets of where Turing Award winners born before 1950 and after 1940 lived, the living places of Turing Award winner Hinton are probably exposed, although the living places may have been deleted in the training due to the privacy concerns. In this work, inspired by the privacy protection in graph embeddings, we propose a privacy-preserving neural graph database (P-NGDB) to alleviate the risks of privacy leakage in NGDBs. We introduce adversarial training techniques in the training stage to force the NGDBs to generate indistinguishable answers when queried with private information, enhancing the difficulty of inferring sensitive information through combinations of multiple innocuous queries. Extensive experiment results on three datasets show that P-NGDB can effectively protect private information in the graph database while delivering high-quality public answers responses to queries.
privacy preserving, knowledge graph, complex query answering
@@ 1. Introduction
Graph databases (graph DBs) play a crucial role in storing, organizing, and retrieving structured relational information and are widely applied in downstream applications, such as recommender systems (Dong, 2018), fraud detection (Pourhabibi et al., 2020), etc. Unfortunately, traditional graph databases often suffer from the limitation of graph incompleteness, which is a prevalent issue in real-world knowledge graphs, like Freebase (Mintz et al., 2009; Bollacker et al., 2008). This incompleteness leads to the exclusion of relevant results, as the graph database may not capture all the necessary relationships and connections between entities by traversing (Bordes et al., 2013; Galárraga et al., 2013).
To solve the drawbacks, extending the concept of Graph DBs, Ren et al. propose neural graph databases (NGDBs) which combine the flexibility of graph data models with the computational capabilities of neural networks, enabling efficient representation, storage, and analysis of interconnected data (Ren et al., 2023).
@@
Figure 1. Privacy risks of Neural Graph Databases (NGDBs) facing malicious queries. NGDBs can retrieve answers given complex logical queries on incomplete graphs (Ren et al., 2023). However, the privacy problem arises due to the storage of graph information in embedding space and the presence of a query engine that facilitates complex logical query answering. To illustrate the issue, consider an example where an attacker attempts to infer private information about Hinton’s living place in the NGDBs. Direct querying privacy can be easily detected by privacy risk detection, however, attackers can leverage the combination of multiple queries to retrieve desired privacy.
@@
As shown in figure 1, NGDBs are designed with two modules: neural graph storage providing unified storage for diverse entities in an embedding space and neural query engine searching answers to input complex queries from the unified storage. By utilizing graph-based neural network techniques, these databases empower expressive and intelligent querying, inference, and knowledge discovery. This capability, known as complex query answering (CQA), aims to identify answers that satisfy given logical expressions (Hamilton et al., 2018; Ren et al., 2020). Logical expressions are often defined in predicate logic forms with relation projection operations, existential quantifiers $\exists$, logical conjunctions $\wedge$, disjunctions $\vee$, etc. The logical forms of $q$ include terms like $Win(V,Turing\ Award)$ and logical operators like conjunction $\wedge$, representing that find where the Turing Award winner who was born in 1960 lived. Query encoding methods are commonly used in Complex Query Answering (CQA). These methods involve parameterizing entities, relations, and logical operators, encoding both queries and entities into the same embedding space simultaneously, and retrieving answers according to the similarity between the query embedding and entity embedding. For example, GQE (Hamilton et al., 2018) and query2box (Ren et al., 2020) encode queries as vector and hyper-rectangle, respectively.
@@
While NGDBs have demonstrated significant achievements in various domains, they do face unique privacy challenges in comparison to traditional graph DBs. One notable risk arises from their generalization ability. Although generalization efficiently handles incomplete knowledge graphs and enriches the retrieval information, it also enables attackers to infer sensitive information from NGDBs by leveraging the composition of multiple complex queries. Figure 1 demonstrates an example of privacy attacks via malicious complex queries: the living place information of an individual, in this case, Hinton as an example, is considered private. Even though Hinton’s residence is omitted during the construction of the knowledge graph, or if direct privacy queries are restricted, a malicious attacker can still infer this sensitive information without directly querying Hinton’s living place. For example, by querying NGDBs with where Turing Award winners born before 1950 and after 1940 lived, where LeCun’s collaborator lived, etc. The intersection of these queries still presents a significant likelihood of exposing the living place of Turing Award winner Hinton.
@@
Graph privacy has consistently been a significant concern in the field of graph research. Multiple studies demonstrate the vulnerability of graph neural networks to various privacy leakage issues. For instance, graph embeddings are vulnerable to attribute inference attacks (Duddu et al., 2020; Gong and Liu, 2018; Zhang et al., 2022), link prediction attacks (He et al., 2021b; Wu et al., 2022), etc. and various protection methods are proposed (Hu and Song, 2023; Hu et al., 2022; Wang et al., 2021). However, these works focus on graph representation learning and have the assumption that the graph embeddings can be reached by the attackers which are not typical scenarios in NGDBs. NGDBs often provide CQA service and do not publish the learned embeddings, and attackers only can infer privacy using thoughtfully designed queries. Furthermore, these works solely study node classification and link prediction problems, leaving the query answering privacy in NGDBs unexplored.
@@
In this paper, we expose the potential privacy risks of NGDBs with formal definition and evaluation. We introduce privacy protection objectives to the existing NGDBs, categorizing the answers of knowledge graph complex queries into private and public domains. In order to safeguard the private information within the knowledge graph, NGDBs should strive to maintain high-quality retrieval of non-private answers for queries, while intentionally obfuscating the private-threatening answers to innocuous yet private queries. As shown in figure 1, the answer sets of queries proposed by privacy attackers are obfuscated, and attackers will face greater challenges in predicting private information, thereby effectively safeguarding the privacy of NGDBs. Meanwhile, we also create the corresponding benchmark on three datasets (Freebase (Bollacker et al., 2008), YAGO (Suchanek et al., 2007), and DBpedia (Bizer et al., 2009)) to evaluate the performance of CQA ability on public query answers and the protection of private information.
@@
To alleviate the privacy leakage problem in NGDBs, we propose Privacy-preserving Neural Graph Databases (P-NGDBs) as a solution. P-NGDBs divide the information in graph databases into private and public parts and categorize complex queries’ answers to various privacy risks based on whether or not they involve sensitive information. P-NGDBs can provide answers with different levels of precision in response to queries with varying privacy risks. We introduce adversarial techniques in the training stage of P-NGDBs to generate indistinguishable answers when queried with private information, enhancing the difficulty of inferring privacy through complex private queries. We summarize our major contributions as follows:
@@
  • We introduce P-NGDB, a privacy protection framework for safeguarding privacy in NGDBs. Extensive experiments conducted on three datasets demonstrate its ability to effectively preserve high information retrieval performance while ensuring privacy protection.
@@ 2. Related Work
In this section, we summarize the related work. Our work is related to complex query answering (CQA) and graph privacy.
@@ 2.1. Complex Query Answering
One essential function of neural graph databases is to answer complex structured logical queries according to the data stored in the NGDBs, namely complex query answering (CQA). Query embedding is one of the prevalent methods for complex query answering because of its effectiveness and efficiency. These methods for query embedding utilize diverse structures to represent complex logical queries and effectively model queries across various scopes. Some methods encode queries into various geometric representations: The GQE (Hamilton et al., 2018) model encodes queries to vectors in embedding space, Query2box (Ren et al., 2020) encodes queries to hyper-rectangle embeddings, HypeE (Liu et al., 2021) encodes queries as hyperbolic embeddings to further support disjunction logical operators; Query2Particle (Bai et al., 2022), ConE (Zhang et al., 2021b) supports negation operators using multiple vector embeddings and cone embeddings, respectively. NRN (Bai et al., 2023a) is proposed to encode numerical values. Besides, some methods try to use probabilistic structures to encode logic queries: BetaE (Ren and Leskovec, 2020), GammaE (Yang et al., 2022), PERM (Choudhary et al., 2021) propose to use Beta, Gamma, and Gaussian distributions to encode logic knowledge graph queries, respectively. Meanwhile, Neural structures are utilized to encode complex queries: BiQE (Kotnis et al., 2021), SQE (Bai et al., 2023b), QE-GNN (Zhu et al., 2022) are proposed to use transformer, sequential encoders, and message-passing graph neural networks to encode queries, respectively. Query decomposition (Arakelyan et al., 2020) proposes to decompose complex queries and LMPNN (Wang et al., 2022) proposes to use one-hop message passing on query graphs to conduct CQA.
@@ 2.2. Graph Privacy
Various privacy protection methods have been proposed to preserve privacy in graph representation models. Anonymization techniques (Liu and Terzi, 2008; Rossi et al., 2015; Hoang et al., 2020; Zheleva and Getoor, 2007) are applied in graphs to reduce the probability of individual and link privacy leakage. Graph summarization (Hay et al., 2008) aims to publish a set of anonymous graphs with private information removed. Learning-based methods are proposed with the development of graph representation learning, (Wang et al., 2021; Liao et al., 2020; Li et al., 2020) regard graph representation learning as the combination of two sub-tasks: primary objective learning and privacy preservation, and use adversarial training to remove the sensitive information while maintaining performance in the original tasks. Meanwhile, some methods (Liu et al., 2022; Oh et al., 2022; Hu and Song, 2023) disentangle the sensitive information from the primary learning objectives. Additionally, differential privacy (Kasiviswanathan et al., 2013; Day et al., 2016; Shen and Yu, 2013) introduces noise into representation models and provides privacy guarantees for the individual privacy of the datasets (Daigavane et al., 2021). Though noise introduced by differential privacy protects models from privacy leakage, it can also impact the performance of the original tasks significantly. Federated learning utilizes differential privacy and encryption to prevent the transmission of participants’ raw data and is widely applied in distributed graph representation learning (Peng et al., 2021; He et al., 2021a). However, federated learning cannot be applied to protect intrinsic private information in the representation models.
@@ 3.1. Preliminary
We consider the privacy problem in the multi-relational knowledge graph with numerical attributes. We denote a knowledge graph as $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{A},\mathcal{N})$, where $\mathcal{V}$ denotes the set of vertices representing entities in the knowledge graph. $\mathcal{R}$ denotes the set of relations where $r\in\mathcal{R}$ is a binary function defined as $r:\mathcal{V}\times\mathcal{V}\rightarrow\{0,1\}$ describing the relation between entities in the knowledge graph, where $r(u,v)=1$ denotes there is a relation $r$ between entities $u$ and $v$ and $r(u,v)=0$ otherwise. $\mathcal{A}$ denotes the set of attributes that can be split into private attributes and public attributes: $\mathcal{A}=\mathcal{A}_{\text{private}}\cup\mathcal{A}_{\text{public}}$, where $a\in\mathcal{A}$ is a binary function defined as $a:\mathcal{V}\times\mathcal{N}\rightarrow\{0,1\}$ describing the numerical attribute of entities, where $a(u,x)=1$ denotes the entity $u$ has the attribute $a$ of value $x\in\mathcal{N}$ and $a(u,x)=0$ otherwise. Besides, there are $f(x_{1},x_{2}):\mathcal{N}\times\mathcal{N}\rightarrow\{0,1\}$ to denote the relation between numerical values. To prevent privacy leakage, the private attribute of entities $\mathcal{A}_{\text{private}}$ cannot be exposed and should be protected from inferences.
The complex query is the key task of NGDBs and can be defined in existential positive first-order logic form, consisting of various types of logic expressions like existential quantifiers $\exists$, logic conjunctions $\wedge$, and disjunctions $\vee$. In the logical expression, there is a unique variable $V_{?}$ in each logic query that denotes our query target. Variables $V_{1},\cdots,V_{k}$ and numerical variables $X_{1},\cdots,X_{l}$ denote existential quantified entities in a complex query. Besides, there are anchor entities $V_{a}$ and values $X_{a}$ with the given content in a query. The complex query aims to identify the target entity $V_{?}$, such that there are $V_{1},\cdots,V_{k}\in\mathcal{V}$ and $X_{1},\cdots,X_{l}\in\mathcal{N}$ in the knowledge graph that can satisfy the given logical expressions. We denote complex query expression in the disjunctive normal form (DNF) in the following:
(1) $\begin{split}q[V_{?}]&=V_{?}.V_{1},...,V_{k},X_{1},...,X_{l}:c_{1}\lor c_{2}\lor...\lor c_{n}\\ c_{i}&=e_{i,1}\land e_{i,2}\land...\land e_{i,m},\end{split}$
where $e_{i,j}$ is the atomic logic expression, which can be $r(V,V^{\prime})$ denotes relation $r$ between entities $V$ and $V^{\prime}$, or $a(V,X)$ denotes attribute $a$ of entities $V$ with value $X$ and $f(X,X^{\prime})$ denotes the relation between numerical attributes. $c_{i}$ is the conjunction of several atomic logic expressions $e_{i,j}$. $V,V^{\prime},X$ are either anchor entities or values, or existentially quantified variables or numearical variables.

### 22 Feb 2024 v3

# Privacy-Preserved Neural Graph Databases

@@ Abstract.
In the era of large language models (LLMs), efficient and accurate data retrieval has become increasingly crucial for the use of domain-specific or private data in the retrieval augmented generation (RAG). Neural graph databases (NGDBs) have emerged as a powerful paradigm that combines the strengths of graph databases (GDBs) and neural networks to enable efficient storage, retrieval, and analysis of graph-structured data which can be adaptively trained with LLMs. The usage of neural embedding storage and Complex neural logical Query Answering (CQA) provides NGDBs with generalization ability. When the graph is incomplete, by extracting latent patterns and representations, neural graph databases can fill gaps in the graph structure, revealing hidden relationships and enabling accurate query answering. Nevertheless, this capability comes with inherent trade-offs, as it introduces additional privacy risks to the domain-specific or private databases. Malicious attackers can infer more sensitive information in the database using well-designed queries such as from the answer sets of where Turing Award winners born before 1950 and after 1940 lived, the living places of Turing Award winner Hinton are probably exposed, although the living places may have been deleted in the training stage due to the privacy concerns. In this work, we propose a privacy-preserved neural graph database (P-NGDB) framework to alleviate the risks of privacy leakage in NGDBs. We introduce adversarial training techniques in the training stage to enforce the NGDBs to generate indistinguishable answers when queried with private information, enhancing the difficulty of inferring sensitive information through combinations of multiple innocuous queries. Extensive experimental results on three datasets show that our framework can effectively protect private information in the graph database while delivering high-quality public answers responses to queries.
Privacy preserving, Knowledge graphs (KGs), Complex query answering (CQA), Neural graph databases (NGDBs)
@@ 1. Introduction
Graph DataBases (GDBs) play a crucial role in storing, organizing, and retrieving structured relational data and support many traditional data-intensive applications, such as recommender systems (Dong, 2018), fraud detection (Pourhabibi et al., 2020), etc. In the era of Large Language Models (LLMs), the role of GDBs gets increasingly important because of the Retrieval Augmented Generation (RAG) paradigm, where the LLM agents can be significantly enhanced by external GDBs such as Knowledge Graphs (KGs). [^1] https://openai.com/research/emergent-tool-use [^2] https://docs.llamaindex.ai/en/stable/examples/agent/openai_agent.html [^3] https://www.langchain.com/use-case/agents Such interaction of LLMs and GDBs enables new possibilities by creating interactive natural language interfaces over structural data for domain-specific applications.
The RAG paradigm of LLMs also highlights new challenges of traditional GDB. The natural language interaction enlarges the possible ways of utilizing the knowledge managed in GDBs, requiring simultaneous knowledge discovery and query answering. Knowledge discovery is particularly important for domain-specific data and knowledge graphs, as the graph database may not capture all the necessary relationships and connections between entities by traversing (Bordes et al., 2013; Galárraga et al., 2013).
@@
Figure 1. Privacy risks of NGDBs facing malicious queries. To illustrate the issue, consider an example where an attacker attempts to infer private information about Hinton’s living place in the NGDBs. Direct querying private information can be easily detected by privacy risk detection, however, attackers can leverage well-designed queries to retrieve desired privacy. In this example, the privacy attacker can query “where Turing Award winners born after 1940 lived,” “where ones before 1950 lived,” and “where LeCun’s collaborator lived,” etc. And the query can be more complex as shown in privacy risk queries. The return answer denoted in red may leak private information in NGDBs. The intersection of these queries still presents a significant likelihood of exposing the living place of Turing Award winner Hinton.
@@
To address the challenges, extending the concept of graph DBs, Ren et al. (Ren et al., 2023) propose neural graph databases (NGDBs) which combine the flexibility of graph data models with the computational capabilities of neural networks, enabling effective and efficient representation, storage, and analysis of interconnected data. As shown in Figure 1, NGDBs are designed with two modules: neural graph storage providing unified storage for diverse entries in an embedding space and neural query engine searching answers to input complex queries from the unified storage (Ren et al., 2023). By utilizing graph-based neural network techniques, these databases empower expressive and intelligent querying, inference, and knowledge discovery. This capability, known as complex query answering (CQA) (Arakelyan et al., 2021), aims to identify answers that satisfy given logical expressions (Hamilton et al., 2018). Query encoding methods are commonly used in CQA. These methods involve parameterizing entities, relations, and logical operators, encoding both queries and entities into the same embedding space simultaneously, and retrieving answers according to the similarity between the query (or set) embeddings and answer embeddings. For example, GQE (Hamilton et al., 2018), Q2B (Ren et al., 2020), and Q2P (Bai et al., 2022) encode queries as vectors, a hyper-rectangle, and a set of particles, respectively. Ideally, CQA can be jointly trained with RAG so that the language model agents can better adapt to domain-specific data (Lee et al., 2019; Guu et al., 2020).
@@
While NGDBs and their CQA ability have demonstrated significant achievements in various domains, they do face unique privacy challenges in comparison to traditional graph DBs. One notable risk arises from their generalization ability. Although generalization efficiently handles incomplete knowledge and enriches the retrieved information, it also enables attackers to infer sensitive information from NGDBs by leveraging the composition of multiple complex queries. Figure 1 demonstrates an example of privacy attacks via malicious complex queries: the living place information of an individual. In this case, Hinton as an example, is considered to be private. Even though Hinton’s residence is omitted during the construction of the knowledge graph, or if direct privacy queries are restricted, a malicious attacker can still infer this sensitive information without directly querying Hinton’s living place but conducting well-designed related queries. The answers denoted in red may leak sensitive information by accident.
@@
Graph privacy has consistently been a significant concern in the field of graph research. Multiple studies demonstrate the vulnerability of graph neural networks to various privacy leakage issues. For instance, graph embeddings are vulnerable to attribute inference attacks (Duddu et al., 2020; Gong and Liu, 2018; Zhang et al., 2022) and link prediction attacks (He et al., 2021b; Wu et al., 2022), and various protection methods are proposed (Hu and Song, 2023; Hu et al., 2022; Wang et al., 2021a). However, these works focus on graph representation learning and have the assumption that the graph embeddings can be reached by the attackers which are not typical scenarios in NGDBs. NGDBs often provide CQA service and do not publish the learned embeddings, and attackers can only infer private information using thoughtfully designed queries. Furthermore, these works solely study node classification and link prediction problems, leaving the query answering privacy in NGDBs unexplored.
@@
In this paper, we expose the potential privacy risks of NGDBs with formal definition and evaluation. We introduce privacy protection objectives to the existing NGDBs, categorizing the answers of knowledge graph complex queries into private and public domains. To safeguard private information, NGDBs should strive to maintain high-quality retrieval of non-private answers for queries, while intentionally obfuscating the private-threatening answers to innocuous yet private queries. As shown in Figure 1, the answer sets of queries proposed by privacy attackers are obfuscated, and attackers will face greater challenges in predicting private information, thereby effectively safeguarding the privacy of NGDBs. Meanwhile, we also create the corresponding benchmark on three datasets (Freebase (Bollacker et al., 2008), YAGO (Suchanek et al., 2007), and DBpedia (Bizer et al., 2009)) to evaluate the performance of CQA on public query answers and the protection of private information.
@@
To alleviate the privacy leakage problem in NGDBs, we propose Privacy-preserved Neural Graph Databases (P-NGDBs) as a solution. P-NGDBs divide the information in graph databases into private and public parts and categorize complex queries’ answers to various privacy risks based on whether or not they involve sensitive information. P-NGDBs can provide answers with different levels of precision in response to queries with varying privacy risks. We introduce adversarial techniques in the training stage of P-NGDBs to generate indistinguishable answers when queried with private information, enhancing the difficulty of inferring privacy through complex private queries. We summarize our major contributions as follows:
@@
  • We introduce P-NGDB, a privacy protection framework for safeguarding privacy in NGDBs. Extensive experiments conducted on three datasets demonstrate its effectiveness in preserving retrieval performance and protecting privacy.
@@ 2.1. Complex Query Answering
One essential function of neural graph databases is to answer complex structured logical queries according to the data stored in the NGDBs, namely complex query answering (CQA) (Wang et al., 2022; Ren et al., 2023). Query encoding is one of the prevalent methods for complex query answering because of its effectiveness and efficiency. These methods for query embedding utilize diverse structures to represent complex logical queries and effectively model queries across various scopes. Some methods encode queries as various geometric structures: such as vectors (Hamilton et al., 2018; Bai et al., 2022), boxes (Ren et al., 2020; Liu et al., 2021), cones (Zhang et al., 2021b) and cones in hyperbolic spaces (Liu et al., 2023). Other methods leverage the probability distributions such as Beta (Ren and Leskovec, 2020), Gaussian (Choudhary et al., 2021), and Gamma (Yang et al., 2022) distributions to encode logic knowledge graph queries. Meanwhile, some approaches motivate the embedding space from the fuzzy logic theory (Chen et al., 2022; Wang et al., 2023a). Advanced neural structures are also utilized to encode complex queries: such as transformers and sequential encoders (Kotnis et al., 2021; Bai et al., 2023c), graph neural networks (Zhu et al., 2022). Pretrained shallow neural link predictors are also adapted into search algorithms (Arakelyan et al., 2021; Bai et al., 2023b; Yin et al., 2024), and learning-to-search frameworks (Wang et al., 2023b; Xu et al., 2023). Notably, NRN (Bai et al., 2023a) is proposed to handle CQA with numerical values, which is of particular interest in privacy protection.
@@ 2.2. Graph Privacy
Various privacy protection methods have been proposed to preserve privacy in graph representation models. Anonymization techniques (Liu and Terzi, 2008; Rossi et al., 2015; Hoang et al., 2020; Zheleva and Getoor, 2007) are applied in graphs to reduce the probability of individual and link privacy leakage. Graph summarization (Hay et al., 2008) aims to publish a set of anonymous graphs with private information removed. Learning-based methods are proposed with the development of graph representation learning, (Wang et al., 2021a; Liao et al., 2020; Li et al., 2020) regard graph representation learning as the combination of two sub-tasks: primary objective learning and privacy preservation, and use adversarial training to remove the sensitive information while maintaining performance in the original tasks. Meanwhile, some methods (Liu et al., 2022; Oh et al., 2022; Hu and Song, 2023) disentangle the sensitive information from the primary learning objectives. Additionally, differential privacy (Kasiviswanathan et al., 2013; Day et al., 2016; Shen and Yu, 2013) introduces noise into representation models and provides privacy guarantees for the individual privacy of the datasets (Daigavane et al., 2021). Though noise introduced by differential privacy protects models from privacy leakage, it can also impact the performance of the original tasks significantly. Federated learning utilizes differential privacy and encryption to prevent the transmission of participants’ raw data and is widely applied in distributed graph representation learning (Peng et al., 2021; He et al., 2021a). However, federated learning cannot be applied to protect intrinsic private information in the representation models.
@@ 3.1. Preliminary
We denote a knowledge graph as $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{A})$, where $\mathcal{V}$ denotes the set of vertices representing entities and attribute values in the knowledge graph. $\mathcal{R}$ denotes the set of relations in the knowledge graph. $\mathcal{A}$ denotes the set of attributes that can be split into private attributes and public attributes: $\mathcal{A}=\mathcal{A}_{\text{private}}\cup\mathcal{A}_{\text{public}}$, where $a\in\mathcal{A}$ is a binary function defined as $a:\mathcal{V}\times\mathcal{V}\rightarrow\{0,1\}$ describing the attribute of entities, where $a(u,x)=1$ denotes the entity $u$ has the attribute $a$ of value $x\in\mathcal{V}$ and $a(u,x)=0$ otherwise. To prevent privacy leakage, the private attribute of entities, we denote the $\mathcal{A}_{\text{private}}$ cannot be exposed and should be protected from inferences.
The complex query is the key task of NGDBs and can be defined in existential positive first-order logic form, consisting of various types of logic expressions like existential quantifiers $\exists$, logic conjunctions $\wedge$, and disjunctions $\vee$. In the logical expression, there is a unique variable $V_{?}$ in each logic query that denotes our query target. Variables $V_{1},\cdots,V_{k}$ denote existentially quantified entities in a complex query. Besides, there are anchor entities $V_{a}$ and values $X_{a}$ with the given content in a query. The complex query aims to identify the target entity $V_{?}$, such that there are $V_{1},\cdots,V_{k}\in\mathcal{V}$ in the knowledge graph that can satisfy the given logical expressions. We denote complex query expression in the disjunctive normal form (DNF) in the following:
(1) $\begin{split}q[V_{?}]&=V_{?}.V_{1},...,V_{k}:c_{1}\lor c_{2}\lor...\lor c_{n}\\ c_{i}&=e_{i,1}\land e_{i,2}\land...\land e_{i,m},\end{split}$
where $e_{i,j}$ is the atomic logic expression, which can be $r(V,V^{\prime})$ denotes relation $r$ between entities $V$ and $V^{\prime}$, and $a(V,V^{\prime})$ denotes attribute $a$ of entities $V$ with value $V^{\prime}$. $c_{i}$ is the conjunction of several atomic logic expressions $e_{i,j}$. $V,V^{\prime}$ are either anchor entities or attributes, or existentially quantified variables.